OpenStudy (anonymous):

Can somebody give me the formula for the Interests principal rate and time i forgot it

OpenStudy (anonymous):

I=PRT for simple interest

OpenStudy (anonymous):

what about if i wanted to get others like rate or principal etc

OpenStudy (anonymous):

You mean how would you get rate if you have interest, principle, and time? or principle if you have the other three?

OpenStudy (anonymous):

both ways please

OpenStudy (anonymous):

You use algebra to get whichever varible you're looking for alone. In order to "undo" the multiplication, you have to divide, right? So if you have I=PRT and you want to find P, you divide by RT, so P=I/(RT). Similarly, if you want to find R, you divide by PT, so R=I/(PT).

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Can you tell me what this means A=(P1+rt)

OpenStudy (anonymous):

Do you mean A=P(1+RT)?

OpenStudy (anonymous):

If that's the equation you're working with, you can distribute the P across to get A=P+PRT, that's telling you that the amount is equal to P (the principle, or amount you started with), plus PRT (which we just found out was the interest.